# Test-time Adaptation for Regression by Subspace Alignment

### v1

@@ D.3 Effect of Original Feature Dimensions on the Subspace
As mentioned in Sec. 3.1, many of the feature dimensions have a small effect on the subspace, which makes the naive feature alignment ineffective. To verify the effect of changing the original features on the subspace, we computed the gradient of a subspace feature vector $\tilde{\mathbf{z}}$ with respect to the $d$-th dimension of the original feature $z_{d}$. With Eq. 7, the norm of the gradient $s_{d}$ is:
@@
Fig. 5 shows the histograms of $s_{d}$ computed with the three datasets. As expected, most of the dimensions of the original feature space had small effects; only a few dimensions had significant effect to the subspace. Specifically, SVHN-MNIST and Biwi Kinect strongly showed this tendency. In contrast, a larger number of raw feature dimensions affected the subspace in UTKFace compared with the other datasets. This is why the test $R^{2}$ score was improved from Source without subspace detection in UTKFace, while the scores on the other datasets were lower than those of Source, as shown in Tab. 6.
@@
Figure 5: Histograms of the gradient of a projected feature $\tilde{z}$ with respect to each original feature dimension $z_{d}$ computed with Eq. 17.
@@ D.4 Additional Ablation
Table 14: Test $R^{2}$ scores of SSA without subspace detection. The bottom row represents the scores with subspace detection of $K=100$ from Tab. 8.
@@
Tab. 12 shows the $R^{2}$ scores on California Housing with changing $K$ of SSA (the same experiment as Tab. 8). $K=10$ produced the best result, but $K=5$ and $25$ also gave competitive scores. When $K\geq 50$, which exceeds the number of the subspace dimensions in Tab. 1, the loss became unstable or diverged because SSA attempted to align too many degenerated feature dimensions, as mentioned in Sec. 4.3.3.
Tabs. 14 and 13 shows the test $R^{2}$ scores without subspace detection (i.e., naively aligning raw features) with respect to $K$. Without subspace detection, the performance of regression on the target domain is limited even when all dimensions of the original feature space are aligned. This result supports the importance of aligning the representative feature subspace with subspace detection.
@@ D.5 Experiments on Vision Transformer
We experimented TTA with vision transformer (ViT) [13] on SVHN-MNIST.
First, we trained a ViT-B/16 regressor on SVHN and computed the number of feature dimensions as Sec. 4.3.4. We used the output of the penultimate layer for the features.
Tab. 15 shows the numbers of the valid (zero variance) and subspace dimensions. Since no activation function is applied to the feature vectors of ViTs unlike ResNets, all 768 feature dimensions are valid (non-zero variance). However, the number of the subspace dimensions is 503 in the regression model, which is smaller that of the classification model. Although the subspace dimension is larger than the ResNet cases in Tab. 1, we can also see the same tendency.
Tab. 16 shows the TTA results on SVHN-MNIST. SSA outperforms the baselines.
@@ D.6 Multi-task Regression Model
We applied TTA methods to multi-task regression models, which output multiple prediction values at the same time. We trained ResNet-50 regression models on Biwi Kinect. Unlike the experiments on the same dataset in Sec. 4.3, a single regression model outputs pitch, roll, and yaw angles.
Tab. 17 shows the test $R^{2}$ score on each target angle. We compared TTA methods that can be easily extended to multi-task regression tasks. Our SSA outperformed the other baselines in most cases. The numbers of the source feature subspace dimensions were 134 and 132 for the male and female models, which are larger than single-task models but much smaller than the entire feature dimensions (2048). Thus, SSA is also effective for multi-task regression models.
Table 17: Test $R^{2}$ scores on multi-task Biwi Kinect. The best scores are in boldface.
@@ D.7 Combination with Classification TTA
Although our SSA is designed for regression, the subspace detection and feature alignment loss can be used to classification since they operates in the feature space and works regardless of model output forms. Here, we applied the subspace detection and feature alignment loss to classification models. We trained ResNet-26 on SVHN and CIFAR10 [32], and tested on MNIST and CIFAR10-C [25].
Tab. 18 shows the number of the valid feature dimensions and subspace dimensions in classification and regression models on the source datasets. Although the number of the feature dimensions of the classification models are larger than those of the regression models, the subspace is smaller than the entire feature space. Thus, we expect that our SSA is also effective on classification models.
Tabs. 19 and 20 show the classification accuracies. We can easily combine our method with entropy-based TTA methods in classification. Although SSA significantly improves the accuracy, combining SSA with Tent [54] further boosts the accuracy.

### v2

@@ D.3 Effect of Original Feature Dimensions on the Subspace
As mentioned in Section 3.1, many of the feature dimensions have a small effect on the subspace, which makes the naive feature alignment ineffective. To verify the effect of changing the original features on the subspace, we computed the gradient of a subspace feature vector $\tilde{\mathbf{z}}$ with respect to the $d$-th dimension of the original feature $z_{d}$. With Equation 6, the norm of the gradient $s_{d}$ is:
@@
Figure 6 shows the histograms of $s_{d}$ computed with the three datasets. As expected, most of the dimensions of the original feature space had small effects; only a few dimensions had significant effect to the subspace. Specifically, SVHN-MNIST and Biwi Kinect strongly showed this tendency. In contrast, a larger number of raw feature dimensions affected the subspace in UTKFace compared with the other datasets. This is why the test $R^{2}$ score was improved from Source without subspace detection in UTKFace, while the scores on the other datasets were lower than those of Source, as shown in Table 6.
@@
Figure 6: Histograms of the gradient of a projected feature $\tilde{z}$ with respect to each original feature dimension $z_{d}$ computed with Equation 15.
@@ D.4 Additional Ablation
Table 14: Test $R^{2}$ scores of SSA without subspace detection. The bottom row represents the scores with subspace detection of $K=100$ from Table 8.
@@
Table 12 shows the $R^{2}$ scores on California Housing with changing $K$ of SSA (the same experiment as Table 8). $K=10$ produced the best result, but $K=5$ and $25$ also gave competitive scores. When $K\geq 50$, which exceeds the number of the subspace dimensions in Table 1, the loss became unstable or diverged because SSA attempted to align too many degenerated feature dimensions, as mentioned in Section 4.3.3.
Tables 14 and 13 shows the test $R^{2}$ scores without subspace detection (i.e., naively aligning raw features) with respect to $K$. Without subspace detection, the performance of regression on the target domain is limited even when all dimensions of the original feature space are aligned. This result supports the importance of aligning the representative feature subspace with subspace detection.
@@ D.5 Experiments on Vision Transformer
We experimented TTA with vision transformer (ViT) (Dosovitskiy et al., 2021) on SVHN-MNIST.
First, we trained a ViT-B/16 regressor on SVHN and computed the number of feature dimensions as Section 4.3.4. We used the output of the penultimate layer for the features.
Table 15 shows the numbers of the valid (zero variance) and subspace dimensions. Since no activation function is applied to the feature vectors of ViTs unlike ResNets, all 768 feature dimensions are valid (non-zero variance). However, the number of the subspace dimensions is 503 in the regression model, which is smaller that of the classification model. Although the subspace dimension is larger than the ResNet cases in Table 1, we can also see the same tendency.
Table 16 shows the TTA results on SVHN-MNIST. SSA outperforms the baselines.
@@ D.6 Multi-task Regression Model
We applied TTA methods to multi-task regression models, which output multiple prediction values at the same time. We trained ResNet-50 regression models on Biwi Kinect. Unlike the experiments on the same dataset in Section 4.3, a single regression model outputs pitch, roll, and yaw angles.
Table 17 shows the test $R^{2}$ score on each target angle. We compared TTA methods that can be easily extended to multi-task regression tasks. Our SSA outperformed the other baselines in most cases. The numbers of the source feature subspace dimensions were 134 and 132 for the male and female models, which are larger than single-task models but much smaller than the entire feature dimensions (2048). Thus, SSA is also effective for multi-task regression models.
Table 17: Test $R^{2}$ scores on multi-task Biwi Kinect (higher is better). The best scores are in boldface.
@@ D.7 Combination with Classification TTA
Although our SSA is designed for regression, the subspace detection and feature alignment loss can be used to classification since they operates in the feature space and works regardless of model output forms. Here, we applied the subspace detection and feature alignment loss to classification models. We trained ResNet-26 on SVHN and CIFAR10 (Krizhevsky et al., 2009), and tested on MNIST and CIFAR10-C (Hendrycks & Dietterich, 2019).
Table 18 shows the number of the valid feature dimensions and subspace dimensions in classification and regression models on the source datasets. Although the number of the feature dimensions of the classification models are larger than those of the regression models, the subspace is smaller than the entire feature space. Thus, we expect that our SSA is also effective on classification models.
Tables 19 and 20 show the classification accuracies. We can easily combine our method with entropy-based TTA methods in classification. Although SSA significantly improves the accuracy, combining SSA with Tent (Wang et al., 2021) further boosts the accuracy.
@@
D.8 Hyperparameter Sensitivity
We investigated SSA’s sensitivity to the other hyperparameters. Tables 21, 22 and 23 and Tables 24, 25 and 26 show the results when varying the learning rate and batch size, respectively.
Typical ranges of the learning rate and batch size produce competitive performance. For the batch size, a larger batch size results in better performance since the estimation of feature mean and variance becomes more accurate. But batch sizes $\geq 16$ produce competitive performance.
Table 21: Test $R^{2}$ scores of SSA on SVHN-MNIST with different learning rates. Higher is better.
Learning rate 0.0001 0.0005 0.001 0.005 0.01
$R^{2}$ ${0.472}_{\pm 0.00}$ ${0.516}_{\pm 0.01}$ ${0.509}_{\pm 0.03}$ ${0.510}_{\pm 0.07}$ ${0.212}_{\pm 0.21}$
Table 22: Test $R^{2}$ scores of SSA on UTKFace with different learning rates. Higher is better.
Learning rate
Defocus blur
Motion blur
Zoom blur
Contrast
Elastic transform
Jpeg compression
Pixelate
Gaussian noise
Impulse noise
Shot noise
Brightness
Fog
Snow
Mean
0.0001 0.7810.781{0.781}0.781 0.8280.828{0.828}0.828 0.8360.836{0.836}0.836 0.7550.755{0.755}0.755 0.8950.895{0.895}0.895 0.8200.820{0.820}0.820 0.9410.941{0.941}0.941 0.5260.526{0.526}0.526 0.5340.534{0.534}0.534 0.4690.469{0.469}0.469 0.8600.860{0.860}0.860 0.4200.420{0.420}0.420 0.4740.474{0.474}0.474 0.7030.703{0.703}0.703
0.0005 0.7860.786{0.786}0.786 0.8280.828{0.828}0.828 0.8340.834{0.834}0.834 0.7670.767{0.767}0.767 0.8930.893{0.893}0.893 0.8200.820{0.820}0.820 0.9380.938{0.938}0.938 0.5510.551{0.551}0.551 0.5660.566{0.566}0.566 0.5130.513{0.513}0.513 0.8610.861{0.861}0.861 0.4340.434{0.434}0.434 0.5080.508{0.508}0.508 0.7150.715{0.715}0.715
0.001 0.7910.791{0.791}0.791 0.8300.830{0.830}0.830 0.8350.835{0.835}0.835 0.7690.769{0.769}0.769 0.8910.891{0.891}0.891 0.8220.822{0.822}0.822 0.9360.936{0.936}0.936 0.5690.569{0.569}0.569 0.5850.585{0.585}0.585 0.5400.540{0.540}0.540 0.8610.861{0.861}0.861 0.4460.446{0.446}0.446 0.5100.510{0.510}0.510 0.7220.722{0.722}0.722
0.005 0.7960.796{0.796}0.796 0.8320.832{0.832}0.832 0.8150.815{0.815}0.815 0.7230.723{0.723}0.723 0.8770.877{0.877}0.877 0.8150.815{0.815}0.815 0.9180.918{0.918}0.918 0.5870.587{0.587}0.587 0.5940.594{0.594}0.594 0.5500.550{0.550}0.550 0.8440.844{0.844}0.844 0.4240.424{0.424}0.424 0.4060.406{0.406}0.406 0.7060.706{0.706}0.706
0.01 0.7690.769{0.769}0.769 0.8040.804{0.804}0.804 0.7690.769{0.769}0.769 0.6680.668{0.668}0.668 0.8630.863{0.863}0.863 0.7920.792{0.792}0.792 0.8990.899{0.899}0.899 0.5230.523{0.523}0.523 0.5740.574{0.574}0.574 0.4640.464{0.464}0.464 0.8110.811{0.811}0.811 0.3270.327{0.327}0.327 0.2930.293{0.293}0.293 0.6580.658{0.658}0.658
Table 23: Test $R^{2}$ scores of SSA on Biwi Kinect with different learning rates. Higher is better.
Female $\rightarrow$ Male Male $\rightarrow$ Female
Learning rate Pitch Roll Yaw Pitch Roll Yaw Mean
0.0001 0.7870.787{0.787}0.787 0.9550.955{0.955}0.955 0.5050.505{0.505}0.505 0.8420.842{0.842}0.842 0.8490.849{0.849}0.849 0.5810.581{0.581}0.581 0.7530.753{0.753}0.753
0.0005 0.8400.840{0.840}0.840 0.9580.958{0.958}0.958 0.5190.519{0.519}0.519 0.8610.861{0.861}0.861 0.8750.875{0.875}0.875 0.5770.577{0.577}0.577 0.7720.772{0.772}0.772
0.001 0.8590.859{0.859}0.859 0.9620.962{0.962}0.962 0.5150.515{0.515}0.515 0.8690.869{0.869}0.869 0.8890.889{0.889}0.889 0.5710.571{0.571}0.571 0.7770.777{0.777}0.777
0.005 0.8770.877{0.877}0.877 0.9630.963{0.963}0.963 0.4920.492{0.492}0.492 0.8590.859{0.859}0.859 0.8930.893{0.893}0.893 0.5490.549{0.549}0.549 0.7720.772{0.772}0.772
0.01 0.8790.879{0.879}0.879 0.9600.960{0.960}0.960 0.4840.484{0.484}0.484 0.8490.849{0.849}0.849 0.8700.870{0.870}0.870 0.4910.491{0.491}0.491 0.7560.756{0.756}0.756
Table 24: Test $R^{2}$ scores of SSA on SVHN-MNIST with different batch sizes. Higher is better
Batch size 8 16 32 64 128 256
$R^{2}$ ${0.353}_{\pm 0.04}$ ${0.497}_{\pm 0.01}$ ${0.505}_{\pm 0.02}$ ${0.509}_{\pm 0.03}$ ${0.528}_{\pm 0.02}$ ${0.522}_{\pm 0.01}$
Table 25: Test $R^{2}$ scores of SSA on UTKFace with different batch sizes. Higher is better.
Batch size
Defocus blur
Motion blur
Zoom blur
Contrast
Elastic transform
Jpeg compression
Pixelate
Gaussian noise
Impulse noise
Shot noise
Brightness
Fog
Snow
Mean
8 0.7790.779{0.779}0.779 0.8040.804{0.804}0.804 0.7960.796{0.796}0.796 0.6510.651{0.651}0.651 0.8670.867{0.867}0.867 0.8040.804{0.804}0.804 0.9140.914{0.914}0.914 0.4990.499{0.499}0.499 0.5160.516{0.516}0.516 0.4160.416{0.416}0.416 0.8340.834{0.834}0.834 0.4080.408{0.408}0.408 0.3900.390{0.390}0.390 0.6680.668{0.668}0.668
16 0.7820.782{0.782}0.782 0.8090.809{0.809}0.809 0.8080.808{0.808}0.808 0.6990.699{0.699}0.699 0.8740.874{0.874}0.874 0.8070.807{0.807}0.807 0.9150.915{0.915}0.915 0.5450.545{0.545}0.545 0.5620.562{0.562}0.562 0.4760.476{0.476}0.476 0.8480.848{0.848}0.848 0.4210.421{0.421}0.421 0.4240.424{0.424}0.424 0.6900.690{0.690}0.690
32 0.7920.792{0.792}0.792 0.8240.824{0.824}0.824 0.8300.830{0.830}0.830 0.7490.749{0.749}0.749 0.8850.885{0.885}0.885 0.8160.816{0.816}0.816 0.9290.929{0.929}0.929 0.5660.566{0.566}0.566 0.5830.583{0.583}0.583 0.5330.533{0.533}0.533 0.8580.858{0.858}0.858 0.4460.446{0.446}0.446 0.4820.482{0.482}0.482 0.7150.715{0.715}0.715
64 0.7910.791{0.791}0.791 0.8300.830{0.830}0.830 0.8350.835{0.835}0.835 0.7690.769{0.769}0.769 0.8910.891{0.891}0.891 0.8220.822{0.822}0.822 0.9360.936{0.936}0.936 0.5690.569{0.569}0.569 0.5910.591{0.591}0.591 0.5400.540{0.540}0.540 0.8610.861{0.861}0.861 0.4460.446{0.446}0.446 0.5100.510{0.510}0.510 0.7220.722{0.722}0.722
128 0.7880.788{0.788}0.788 0.8280.828{0.828}0.828 0.8350.835{0.835}0.835 0.7800.780{0.780}0.780 0.8930.893{0.893}0.893 0.8220.822{0.822}0.822 0.9390.939{0.939}0.939 0.5810.581{0.581}0.581 0.5900.590{0.590}0.590 0.5450.545{0.545}0.545 0.8620.862{0.862}0.862 0.4450.445{0.445}0.445 0.5230.523{0.523}0.523 0.7250.725{0.725}0.725
256 0.7890.789{0.789}0.789 0.8230.823{0.823}0.823 0.8360.836{0.836}0.836 0.7700.770{0.770}0.770 0.8960.896{0.896}0.896 0.8210.821{0.821}0.821 0.9400.940{0.940}0.940 0.5400.540{0.540}0.540 0.5940.594{0.594}0.594 0.5510.551{0.551}0.551 0.8700.870{0.870}0.870 0.4390.439{0.439}0.439 0.5220.522{0.522}0.522 0.7230.723{0.723}0.723
Table 26: Test $R^{2}$ scores of SSA on Biwi Kinect with different batch sizes. Higher is better.
Female $\rightarrow$ Male Male $\rightarrow$ Female
Batch size Pitch Roll Yaw Pitch Roll Yaw Mean
8 0.8720.872{0.872}0.872 0.9300.930{0.930}0.930 0.4500.450{0.450}0.450 0.8190.819{0.819}0.819 0.7700.770{0.770}0.770 0.4680.468{0.468}0.468 0.7180.718{0.718}0.718
16 0.8700.870{0.870}0.870 0.9530.953{0.953}0.953 0.4630.463{0.463}0.463 0.8580.858{0.858}0.858 0.8570.857{0.857}0.857 0.5250.525{0.525}0.525 0.7540.754{0.754}0.754
32 0.8680.868{0.868}0.868 0.9610.961{0.961}0.961 0.4980.498{0.498}0.498 0.8680.868{0.868}0.868 0.8860.886{0.886}0.886 0.5620.562{0.562}0.562 0.7740.774{0.774}0.774
64 0.8590.859{0.859}0.859 0.9620.962{0.962}0.962 0.5150.515{0.515}0.515 0.8690.869{0.869}0.869 0.8890.889{0.889}0.889 0.5710.571{0.571}0.571 0.7770.777{0.777}0.777
128 0.8410.841{0.841}0.841 0.9600.960{0.960}0.960 0.5240.524{0.524}0.524 0.8630.863{0.863}0.863 0.8830.883{0.883}0.883 0.5690.569{0.569}0.569 0.7730.773{0.773}0.773
256 0.8190.819{0.819}0.819 0.9570.957{0.957}0.957 0.5210.521{0.521}0.521 0.8550.855{0.855}0.855 0.8700.870{0.870}0.870 0.5830.583{0.583}0.583 0.7680.768{0.768}0.768
D.9 Additional Results
Tables 27 and 28 provide the performance measured by MAE on UTKFace and Biwi Kinect, which are corresponding to Tables 4 and 5.
Table 27: Test MAE scores on UTKFace with image corruption (lower is better). The best scores are bolded.
Method
Defocus blur
Motion blur
Zoom blur
Contrast
Elastic transform
Jpeg compression
Pixelate
Gaussian noise
Impulse noise
Shot noise
Brightness
Fog
Snow
Mean
Source $10.12$ $12.64$ $7.68$ $39.99$ $7.92$ $13.91$ $9.65$ $32.00$ $32.02$ $31.92$ $8.81$ $15.52$ $19.76$ $18.61$
DANN $9.44$ $8.88$ $8.73$ $19.59$ $7.52$ $8.01$ $6.09$ $41.26$ $35.56$ $38.21$ $9.11$ $16.07$ $18.11$ $17.43$
TTT $6.78$ $6.55$ $6.61$ $6.51$ $5.77$ $6.59$ $5.13$ $9.56$ $9.46$ $10.00$ $6.51$ $10.96$ $9.84$ $7.71$
BN-Adapt $7.23$ $6.91$ $6.69$ $7.59$ $5.97$ $6.81$ $5.40$ $10.37$ $10.32$ $10.99$ $6.44$ $11.51$ $10.55$ $8.21$
Prototype $21.52$ $21.62$ $21.65$ $20.00$ $21.28$ $21.02$ $21.27$ $18.46$ $18.45$ $18.42$ $21.71$ $20.71$ $20.70$ $20.52$
FR $6.12$ $\mathbf{5.47}$ 5.155.15{5.15}5.15 $6.29$ $4.47$ $5.85$ $\mathbf{3.15}$ $9.99$ $9.87$ $10.53$ $5.03$ $11.04$ $10.36$ $7.18$
VM $28.28$ $28.14$ $28.42$ $27.45$ $27.75$ $27.74$ $26.97$ $29.07$ $29.21$ $29.05$ $27.80$ $29.50$ $29.19$ $28.35$
RSD $6.35$ $5.68$ $5.25$ $6.60$ $4.61$ $5.93$ $3.34$ $10.26$ $10.35$ $10.94$ $5.12$ $11.13$ $9.98$ $7.35$
SSA (ours) $\mathbf{6.05}$ 5.525.52{5.52}5.52 $\mathbf{5.09}$ $\mathbf{6.05}$ $\mathbf{4.46}$ $\mathbf{5.72}$ 3.273.27{3.27}3.27 $\mathbf{9.05}$ $\mathbf{8.95}$ $\mathbf{9.29}$ $\mathbf{5.01}$ $\mathbf{10.60}$ $\mathbf{9.50}$ $\mathbf{6.81}$
Oracle $5.21$ $4.55$ $4.53$ $4.96$ $3.96$ $4.92$ $2.64$ $8.44$ $8.28$ $8.42$ $4.45$ $10.03$ $8.07$ $6.03$
Table 28: Test MAE scores on Biwi Kinect (lower is better). The best scores are bolded.
Female $\rightarrow$ Male Male $\rightarrow$ Female
Method Pitch Roll Yaw Pitch Roll Yaw Mean
Source $0.150$ $0.081$ $0.087$ $0.160$ $0.171$ $0.093$ $0.124$
DANN $0.163_{\pm 0.01}$ $0.163_{\pm 0.01}$ $0.136_{\pm 0.01}$ $0.181_{\pm 0.01}$ $0.166_{\pm 0.01}$ $0.147_{\pm 0.00}$ $0.159_{\pm 0.00}$
TTT $0.283_{\pm 0.02}$ $0.236_{\pm 0.00}$ $0.140_{\pm 0.00}$ $0.165_{\pm 0.00}$ $0.212_{\pm 0.00}$ $0.190_{\pm 0.00}$ $0.205_{\pm 0.00}$
BN-adapt $0.146_{\pm 0.00}$ $0.085_{\pm 0.00}$ $0.090_{\pm 0.00}$ $0.153_{\pm 0.00}$ $0.134_{\pm 0.00}$ $\mathbf{0.090_{\pm 0.00}}$ $0.116_{\pm 0.00}$
Prototype $6.935_{\pm 0.00}$ $6.935_{\pm 0.00}$
FR $0.376_{\pm 0.05}$ $0.192_{\pm 0.01}$ $0.248_{\pm 0.02}$ $0.212_{\pm 0.01}$ $0.150_{\pm 0.01}$ $0.182_{\pm 0.02}$ $0.226_{\pm 0.01}$
VM $0.367_{\pm 0.00}$ $0.407_{\pm 0.00}$ $0.136_{\pm 0.00}$ $0.418_{\pm 0.00}$ $0.463_{\pm 0.00}$ $0.135_{\pm 0.00}$ $0.321_{\pm 0.00}$
RSD $0.142_{\pm 0.01}$ $0.085_{\pm 0.00}$ $0.090_{\pm 0.00}$ $0.154_{\pm 0.00}$ $0.132_{\pm 0.00}$ $0.121_{\pm 0.00}$
SSA (ours) $\mathbf{0.112_{\pm 0.00}}$ $\mathbf{0.079_{\pm 0.00}}$ $\mathbf{0.086_{\pm 0.00}}$ $\mathbf{0.141_{\pm 0.00}}$ $\mathbf{0.126_{\pm 0.00}}$ $\mathbf{0.090_{\pm 0.00}}$ $\mathbf{0.106_{\pm 0.00}}$
Oracle $0.054_{\pm 0.00}$ $0.054_{\pm 0.00}$ $0.059_{\pm 0.00}$ $0.076_{\pm 0.00}$ $0.068_{\pm 0.00}$ $0.065_{\pm 0.00}$ $0.063_{\pm 0.00}$
D.10 Online Setup
We evaluated regression TTA in an batched online setting, where model update and evaluation are performed alternatively with a batch in every iteration. Tables 29, 30 and 31 display the results. Our SSA can outperform the baselines also in the batched online setting.
Table 29: Test scores on SVHN-MNIST in the batched online setting. The best scores are bolded.
Method $R^{2}(\uparrow)$ RMSE ($\downarrow$) MAE ($\downarrow$)
Source 0.4060.406{0.406}0.406 2.2322.232{2.232}2.232 1.6081.608{1.608}1.608
TTT ${0.296_{\pm 0.01}}$ ${2.430_{\pm 0.01}}$ ${1.587_{\pm 0.01}}$
BN-adapt ${0.384_{\pm 0.00}}$ ${2.272_{\pm 0.00}}$ ${1.480_{\pm 0.00}}$
Prototype ${0.484_{\pm 0.00}}$ ${2.080_{\pm 0.00}}$ ${1.489_{\pm 0.00}}$
FR ${0.342_{\pm 0.00}}$ ${2.348_{\pm 0.00}}$ ${1.657_{\pm 0.01}}$
VM ${-227.105_{\pm 7.22}}$ ${43.729_{\pm 0.69}}$ ${37.021_{\pm 0.72}}$
RSD ${0.312_{\pm 0.08}}$ ${2.397_{\pm 0.15}}$ ${1.607_{\pm 0.15}}$
SSA (ours) $\mathbf{0.488_{\pm 0.01}}$ $\mathbf{2.072_{\pm 0.03}}$ $\mathbf{1.265_{\pm 0.02}}$
Oracle ${0.745_{\pm 0.00}}$ ${1.463_{\pm 0.01}}$ ${0.882_{\pm 0.01}}$
Table 30: Test $R^{2}$ scores on UTKFace with image corruption in the batched online setting. The best scores are bolded.
Method
Defocus blur
Motion blur
Zoom blur
Contrast
Elastic transform
Jpeg compression
Pixelate
Gaussian noise
Impulse noise
Shot noise
Brightness
Fog
Snow
Mean
Source $0.410$ $0.187$ $0.658$ $-3.906$ $0.701$ $0.069$ $0.595$ $-2.536$ $-2.539$ $-2.522$ $0.661$ $-0.018$ $-0.543$ $-0.676$
TTT $0.742$ $0.758$ $0.773$ $0.778$ $0.828$ $0.769$ $0.860$ $0.519$ $0.531$ $0.483$ $0.776$ $0.391$ $0.462$ $0.667$
BN-Adapt $0.726$ $0.758$ $0.757$ $0.719$ $0.822$ $0.774$ $0.849$ $0.492$ $0.504$ $0.451$ $0.788$ $0.375$ $0.437$ $0.650$
Prototype $-1.001$ $-1.017$ $-1.014$ $-0.719$ $-0.965$ $-0.907$ $-0.972$ $-0.514$ $-0.513$ $-0.511$ $-1.003$ $-0.823$ $-0.822$ $-0.829$
FR $0.786$ $\mathbf{0.834}$ $\mathbf{0.844}$ $0.765$ $0.895$ $0.820$ $\mathbf{0.944}$ $0.499$ $0.509$ $0.449$ $0.858$ $0.409$ $0.453$ $0.697$
VM $-1.457$ $-1.401$ $-1.503$ $-1.230$ $-1.327$ $-1.343$ $-1.251$ $-1.538$ $-1.538$ $-1.524$ $-1.330$ $-1.632$ $-1.577$ $-1.435$
RSD $0.783$ $0.829$ $0.843$ $0.761$ $0.893$ $0.820$ $0.940$ $0.503$ $0.509$ $0.447$ $0.858$ $0.416$ $0.483$ $0.699$
SSA (ours) $\mathbf{0.794}$ $\mathbf{0.834}$ $\mathbf{0.844}$ $\mathbf{0.789}$ $\mathbf{0.896}$ $\mathbf{0.823}$ 0.9430.943{0.943}0.943 $\mathbf{0.550}$ $\mathbf{0.564}$ $\mathbf{0.524}$ $\mathbf{0.861}$ $\mathbf{0.427}$ $\mathbf{0.520}$ $\mathbf{0.721}$
Oracle $0.825$ $0.868$ $0.863$ $0.824$ $0.908$ $0.845$ $0.951$ $0.578$ $0.586$ $0.577$ $0.872$ $0.472$ $0.605$ $0.752$
Table 31: Test $R^{2}$ scores on Biwi Kinect in the batched online setting. The best scores are bolded.
Female $\rightarrow$ Male Male $\rightarrow$ Female
Method Pitch Roll Yaw Pitch Roll Yaw Mean
Source $0.759$ $0.956$ $0.481$ $0.763$ $0.791$ $0.485$ $0.706$
TTT $-0.207_{\pm 0.05}$ $0.610_{\pm 0.00}$ $0.010_{\pm 0.02}$ $0.743_{\pm 0.00}$ $0.722_{\pm 0.00}$ $-0.296_{\pm 0.00}$ $0.264_{\pm 0.01}$
BN-adapt $0.759_{\pm 0.00}$ $0.951_{\pm 0.00}$ $0.487_{\pm 0.00}$ $0.827_{\pm 0.00}$ $0.837_{\pm 0.00}$ $\mathbf{0.567_{\pm 0.00}}$ $0.738_{\pm 0.00}$
Prototype $-317.740_{\pm 0.02}$
FR $-0.124_{\pm 0.06}$ $0.818_{\pm 0.02}$ $-2.049_{\pm 0.24}$ $0.775_{\pm 0.02}$ $0.852_{\pm 0.01}$ $0.128_{\pm 0.08}$ $0.067_{\pm 0.06}$
VM $-0.242_{\pm 0.01}$ $-0.057_{\pm 0.00}$ $-0.089_{\pm 0.00}$ $-0.101_{\pm 0.00}$ $-0.051_{\pm 0.00}$ $-0.006_{\pm 0.00}$ $-0.091_{\pm 0.00}$
RSD $0.768_{\pm 0.01}$ $0.951_{\pm 0.00}$ $0.486_{\pm 0.00}$ $0.826_{\pm 0.00}$ $0.838_{\pm 0.00}$
SSA (ours) $\mathbf{0.825_{\pm 0.00}}$ $\mathbf{0.957_{\pm 0.00}}$ $\mathbf{0.502_{\pm 0.00}}$ $\mathbf{0.853_{\pm 0.00}}$ $\mathbf{0.865_{\pm 0.00}}$ $\mathbf{0.567_{\pm 0.00}}$ $\mathbf{0.761_{\pm 0.00}}$
Oracle $0.923_{\pm 0.00}$ $0.971_{\pm 0.00}$ $0.672_{\pm 0.00}$ $0.925_{\pm 0.00}$ $0.934_{\pm 0.00}$ $0.717_{\pm 0.01}$ $0.857_{\pm 0.00}$